OpenStudy (anonymous):

for a medal and a fan!!!!!!!!!! Use any method to solve the equation. If necessary, round to the nearest hundredth. 2x^2 + x = 10

OpenStudy (anonymous):

7.66

OpenStudy (texaschic101):

quadratic formula x = (-b (+-) sq rt b^2 - 4ac) / (2a) 2x^2 + x = 10 2x^2 + x - 10 = 0 a = 2, b = 1, c = -10 x = (-1 (+-) sq rt 1^2 - 4(2)(-10)) / 2(2) x = -1 (+-) sq rt 1 + 80) / 4 x = -1 (+-) sq rt 81) / 4 x = (-1 (+-) 9)/4 x = (-1 + 9)/4 x = 8/4 x = 2 x = (-1 - 4)/2 x = -5/2 so x = 2 and x = -5/2

OpenStudy (texaschic101):

oops.. x = (-1 - 9) / 4 x = -10/4 x = - 5/2 same answer

OpenStudy (texaschic101):

read more here.. http://www.purplemath.com/modules/quadform.htm